What is the ratio between the intensity of fundamental and the third harmonic in a string plucked at the midpoint of length

Grade:12th pass

1 Answers

Arun
6 years ago
 
Dear Ranjit
 
If I1 & I2  are the intensities of the fundamental and third harmonic, respectively then
I1/I3  = 81 [sin2(\pia/L)/sin2(3\pia/L)]
 
Now if the string is plucked at its mid point then a = L/2
hence
I1/I3   = 81
 
Regards
Arun